A conducting cylinder of radius 3.25 cm and length 5.6 cm has a total charge of 5.5 x 10^-9 C distributed uniformly on its surface area. Find for the potential at its surface.

1
Expert's answer
2022-03-01T11:12:44-0500

The potential of electric field

"V=-\frac{\lambda}{2\pi\epsilon_0}\ln r=-\frac{Q}{2\pi\epsilon_0*\pi r^2l}\ln r\\
=-\frac{5.5*10^{-9}}{2\pi*8.85*10^{-12}*\pi (0.0325)^2*0.056}\ln 0.0325\\
=3.5*10^6\:\rm V"